# MOC-RVQ: Multilevel Codebook-assisted Digital Generative Semantic Communication

### v1

@@ III-C Codebook Reordering
As discussed in Section I, a discrepancy exists in the local relationship between code indices and code vectors, rendering VQ-based semantic communication systems susceptible to channel noise. In response to this challenge, we propose a heuristic solution through the codebook reordering (CR) algorithm outlined in Algorithm 1. The algorithm systematically searches for the nearest neighbor vector within $\mathcal{C}$, constructing a new vector sequence, which is then assembled to form a preliminarily sorted codebook. Furthermore, to align with the design principles of Gray code mapping in digital constellation modulation, we also integrate Gray code mapping into our codebook reordering algorithm. The basic idea is to establish proximity relationships under the context of Gray code mapping. For instance, given a 3-bit Gray code sequence represented in decimals as $g=\{0,1,3,2,6,7,5,4\}$, it becomes apparent that "2" and "6" are neighboring elements within the framework of Gray code mapping. Thus, it is essential to allocate two closely related code vectors to these indices.
@@ 8       $\mathcal{C}^{*}[g,:]\leftarrow$ $\mathcal{C}^{*}$ ;
After codebook reordering, the semantic distance between two adjacent Gray code vectors in the updated codebook $\mathcal{C}^{*}$ is reduced compared to the previous disordered version. This addresses the issue where a small jump in the code index could lead to a significant jump in the semantic vector.
@@ IV Experiments
Figure 3: Experimental comparison of methods using PSNR, SSIM, and LPIPS metrics over an additive white Gaussian noise (AWGN) channel with SNR varying from -5 to 30. L1 to L4 denote the level of quantization of our proposed MOC-RVQ, which directly determines the number of bits that need to be transmitted for communication.
@@
Figure 4: Ablation study to comprehend the influence of our proposed noise reduction block (NRB) and codebook reordering (CR) algorithm. Note that all experiments are conducted on the basis of L4 transmission.
IV-A Implementation Details
@@
In accordance with Chen et al. [13], we employ a baseline autoencoder with a downsampling factor $f=8$, indicating that the resolution of latent features is $1/f^{2}$ of the input resolution. To ensure effective model training, we assemble a comprehensive training set by integrating diverse datasets: DIV2K [14], Flickr2K [15], DIV8K [16], and a dataset comprising 10,000 facial images from FFHQ [17]. The generation of training patches involves the following steps: (1) extraction of non-overlapping patches with dimensions $512\times 512$; (2) filtration of patches with minimal texture variation; (3) for the FFHQ dataset, a random resizing operation is applied with scale factors in the range [0.5, 1.0] before cropping. For testing, we uniformly sample 25 high resolution images (with average resolution of $2012\times 1407$) from DIV2K validation set as our test set to simulate modern image transmission scenarios.
@@
We employ the Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$ for both stages, maintaining a learning rate of $0.0001$ throughout the training process. The input image size is randomly cropped to $256\times 256$ with a batch size of 16. The model pretraining stage requires approximately 3 days on 2 GeForce RTX 3090 GPUs, and the NRB finetuning stage also takes about 3 days on the same device.
@@ IV-B Simulations and Discussions
Figure 5: Visualizing the impact of NRB across varying channel conditions. Even in poor channel quality (SNR=5), our full model impressively reconstructs images with clear semantic meaning, demonstrating the robustness of VQ-based semantic communication systems with the assistance of NRB.
In this section, we empirically demonstrate the qualitative and quantitative results of our approach, conduct model ablations, and compare our method with two state-of-the-art baselines. Note that all approaches are under 64-QAM. For channel simulation, initially, we simulate an additive white Gaussian noise (AWGN) channel with SNR varying from -5 to 30 to encompass low, medium, and high-quality channel conditions. Subsequently, we finetune our proposed NRB under this AWGN channel to enhance image reconstruction.
@@
For comparison, we apply the better portable graphics (BPG [18]) and the joint picture expert group (JPEG [19]) as baselines. Notably, both of these image codec formats are susceptible to bit errors, leading to potential image format errors and decoding failures. To mitigate this, we introduce a widely used channel error correction code called Low-Density Parity-Check (LDPC) by setting the LDPC block length and bit rate as 648 bits and 1/2, respectively. Thanks to the self-restoring capability of our proposed method, we do not apply any channel error correction code, which reduces the number of transmitted bits and significantly improves the transmission efficiency of the proposed communication system.
As depicted in Fig. 3, we utilize PSNR, SSIM, and LPIPS as metrics for various methods with a similar number of transmitted symbols over the AWGN channel. LPIPS [20] is specifically introduced for evaluating the perceptual quality of generated images in semantic communication scenarios. Our proposed method exhibits remarkable robustness in deteriorating channel quality, outperforming traditional image coding even without channel error correction mechanism. This is attributed to the VQ-based semantic communication system’s independence from encoding formats, where minor bit changes do not lead to decoding failures. Additionally, the incorporation of NRB and codebook reordering further enhances the robustness of the semantic communication system.
Figure 3 also reveals that lowering the transmission level (L4 to L1) of our model reduces the number of transmitted bits, but it also compromises reconstruction performance. Especially, we can clearly observe the gaps at different levels in LPIPS metrics, which implies that we can dynamically adjust the number of bits that need to be transmitted based on the quality of the channel and the receiver’s requirements for image quality.
In Figure 4, we present an ablation study to understand the impact of our proposed NRB and the CR algorithm. Focusing on LPIPS scores, which better capture perceptual quality, we observe the crucial role of NRB in reconstruction performance, especially under poor channel quality. Simultaneously, the proposed codebook reordering algorithm provides additional gains in image reconstruction performance.
@@
Figure 5 illustrates the visual effect of NRB. With the assistance of NRB, even under relatively poor channel quality (SNR=5), our full model impressively reconstructs images with clear semantic meaning, which greatly expands the working conditions of VQ-based semantic communication systems.
@@ V Conclusion
We introduce MOC-RVQ, a novel two-stage framework addressing challenges in vector quantization and digital constellation modulation for semantic communication systems. By incorporating a multi-head octonary codebook (MOC) to compress the index range and a residual vector quantization (RVQ) mechanism for efficient multilevel semantic communication, our approach demonstrates superior performance over existing methods like BPG and JPEG, even without channel error correction coding. The innovative two-stage model training framework, featuring a noise reduction block (NRB), enhances the system’s ability to restore high-quality semantic features in the presence of channel noise. Additionally, a heuristic codebook reordering algorithm improves system robustness against channel noise, mitigating the inherent mismatch between vector quantization and digital constellation modulation. MOC-RVQ represents a significant advancement in generative semantic communication, providing a promising solution for the joint design challenges and paving the way for further exploration in intelligent communication systems.

### v2

@@ III-C Codebook Reordering
As discussed in Section I, a discrepancy exists in the local relationship between code indices and code vectors. In response to this challenge, before commencing Stage 2, we propose a heuristic solution through a codebook reordering (CR) algorithm outlined in Algorithm 1. The algorithm systematically searches for the nearest neighbor vector within $\mathcal{C}$, resulting in a new vector sequence, which is then assembled to form a preliminarily sorted codebook. Furthermore, to align with the Gray code mapping process in digital constellation modulation, we also integrate Gray code mapping into our CR algorithm (optional). The basic idea is to establish proximity relationships under the context of Gray code mapping. For instance, given a 3-bit Gray code sequence represented in decimals as $g=\{0,1,3,2,6,7,5,4\}$, it becomes apparent that "2" and "6" are neighboring elements. Thus, it is essential to allocate two closely related code vectors to these indices.
1
@@ 11 $\mathcal{C}^{*}[g,:]\leftarrow$ $\mathcal{C}^{*}$ ;
After codebook reordering, the semantic distance between two adjacent Gray code vectors in the updated codebook $\mathcal{C}^{*}$ is reduced compared to the previous disordered version. This addresses the issue where a small jump in the code index would lead to a significant jump in the semantic vector.
@@ IV Experiments
In this section, we empirically demonstrate the qualitative and quantitative results of our model, conduct model ablation, and compare our model with different baselines.
@@
Figure 3: Experimental comparison using PSNR, SSIM, and LPIPS metrics over AWGN channels with SNR from -5 to 30. L1 to L4 correspond to different quantization levels of MOC-RVQ, with L1 as a baseline for the VQGAN-based variant. JSCC-96 is an analog baseline, with a CBR six times higher than L4, and requires analog modulation or full-resolution constellation, which hinders compatibility with current digital communication systems.
@@
Figure 4: Ablation study to comprehend the influence of our proposed noise reduction block (NRB) and codebook reordering (CR) algorithm. Note that all experiments are conducted under L4 transmission, and ’w/o’ denotes ’without’. There are clear gaps between different setting, which verifies the effectiveness of the proposed NRB and CR designs.
IV-A Experimental Setups
@@
Similar to [11], we utilize the autoencoder from VQGAN [10] with a downsampling factor of $f=8$, meaning that the resolution of the bottleneck feature is $1/f^{2}$ of the input, as the backbone. For effective model training, we assemble a comprehensive training set by integrating: DIV2K [12], Flickr2K [13], DIV8K [14], and a dataset comprising 10,000 facial images from FFHQ [15]. Then the training patches are generated by the following steps: (1) Extraction of non-overlapping patches with a resolution of $512\times 512$; (2) Filtration of patches with minimal texture variation; (3) For the FFHQ dataset, a random resizing operation is applied with scale factors in the range [0.5, 1.0] before cropping. For testing, we evenly sample 25 high-resolution images ($2012\times 1407$ in average) from the DIV2K validation set to simulate modern image transmission scenarios.
@@
We employ an Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.99$, and maintain a learning rate of $0.0001$ throughout the training process. The input image size is randomly cropped to $256\times 256$ for efficient training with a batch size of 16. The model pretraining stage requires approximately 3 days on 2 GeForce RTX 3090 GPUs, while the NRB finetuning stage also takes about 3 days on the same device.
Channel Settings and Evaluation Metrics
We simulate AWGN channels with SNR varying from -5 to 30 to encompass low, medium, and high-quality channel conditions. Note that the proposed NRB is finetuned under these AWGN channels and all digital models are equipped 64-QAM modulation. For evaluation, PSNR, SSIM, and LPIPS are utilized as metrics for various methods. LPIPS [16] is specifically introduced for evaluating the perceptual quality of generated images under the semantic communication scenarios.
@@ IV-B Simulations and Discussions
Figure 5: Visualization of removing both the NRB and CR across varying channel conditions. Even in poor channel quality (SNR=0), our full model impressively reconstruct images with clear meaning, demonstrating the robustness of VQ-based semantic communication systems with the assistance of NRB and CR.
@@
We apply the better portable graphics (BPG [17]) and the joint picture expert group (JPEG [18]) as digital baselines by setting the number of bits after compression similar to our proposed MOC-RVQ. Notably, both of these image codec formats are susceptible to bit errors, leading to potential image format errors and decoding failures. To mitigate this, we introduce a widely used channel error correction code called Low-Density Parity-Check (LDPC) by setting the LDPC block length and bit rate as 648 bits and 1/2, respectively. Thanks to the self-restoring capability of our proposed method, we do not apply any channel error correction code for MOR-RVQ.
Additionally, we implement an analog JSCC baseline, referred to as JSCC-96, by retaining only the autoencoder from MOC-RVQ and removing all its digital components. Without loss of generality, we set the number of channels in the bottleneck feature (the encoder output) to 96 for comparison, resulting in a channel bandwidth ratio (CBR) that is six times greater than that of MOC-RVQ-L4. The definition of CBR is as follows:
$CBR=\frac{B}{H\times W\times 3},$ (11)
where $B$ denotes the number of transmitted complex symbols, and $H,W$ represent the width and height of the input RGB image, respectively. Note that the symbols of analog scheme require full-resolution constellation, while the symbols of our method only need few discrete states.
From Fig. 3, we can observe that our proposed method exhibits robustness in deteriorating channel quality when compared with traditional image coding, even in the absence of a channel error correction mechanism. Additionally, while the analog scheme JSCC-96 demonstrates commendable performance, it requires a higher CBR and faces compatibility challenges with existing digital communication systems.
Figure 3 also shows that reducing the transmission level (from L4 to L1) in our model, which directly decreases the number of transmitted bits, leads to a compromise in reconstruction performance. Notably, the gaps at different levels are clearly visible in the LPIPS metrics. It is also worth mentioning that L1 can serve as a special VQGAN-based digital baseline for comparison.
Additionally, We also conduct ablation study (illustrated in Fig. 4) to assess the effects of our proposed NRB and the CR algorithm. We can observe that the proposed NRB plays a crucial role in the reconstruction performance. Simultaneously, the proposed CR algorithm contributes additional gains to the image reconstruction performance.
@@
Figure 5 illustrates the visual effect of removing both the NRB and CR. With the assistance of both these designs, even under relatively poor channel quality (SNR=0), our full model can impressively reconstruct images with clear semantic meaning, thus significantly expanding the working conditions of VQ-based semantic communication systems.
@@ V Conclusion
In this work, we propose MOC-RVQ, a novel approach that combines a multi-head octonary codebook (MOC) with residual vector quantization (RVQ) in a two-stage training framework to address the challenges between vector quantization and digital constellation modulation. By leveraging MOC for codebook compression and RVQ for efficient multilevel semantic communication, our method surpasses traditional techniques like BPG and JPEG, and achieves performance comparable to analog JSCC. Furthermore, the two-stage framework, enhanced by the NRB design and CR algorithm, significantly improves the system’s ability to recover high-quality semantic features under channel noise. While effective, MOC-RVQ is specifically optimized for 64-QAM modulation, suggesting the need for codebook adjustments and retraining for other modulation schemes, providing avenues for future research.